OpenStudy (anonymous):

Find the inverse f(x)=(x+3)/(x-2)

OpenStudy (anonymous):

so first switch x and y

OpenStudy (anonymous):

x=(y+3)/(y-2)

OpenStudy (anonymous):

so after that you can cross multiply to get rid of the denominator

OpenStudy (anonymous):

which gives you x(y-2)=(y+3)

OpenStudy (anonymous):

so now what step do you think you could take to isolate y

OpenStudy (anonymous):

distribute the x(y-2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then you have xy-2x=y+3

OpenStudy (anonymous):

you can add the 2x and then subtract y to get xy-y=2x+3

OpenStudy (anonymous):

now what do you think is the next step?

OpenStudy (anonymous):

@recon14193 my computer lagged, i'm back

OpenStudy (anonymous):

divide by x

OpenStudy (anonymous):

y=(2x+3)/(x-1)

OpenStudy (anonymous):

well you go the right answer but youre not diving by x you are factoring out y so that you have y(x-1)=2x+3

OpenStudy (anonymous):

and then you divide by x-1 to get you had stated before

OpenStudy (anonymous):

thanks